A ramp is pulled out of the back of a truck (on flat ground). There is a 38 degree angle between the ramp and the pavement. If the distance from the base of the ramp to the back of the truck is 10 feet. How long is the ramp? Show sketch and setup. Answer to the nearest tenth of a foot.
Answer by mananth(16946)


.
Cos 38 = BC/AC
Cos 38 = 10/AC
AC = 10/Cos 38
12.7 ft
Ramp length = 12.7 ft